# ConCISE: Confidence-guided Compression in Step-by-step Efficient Reasoning

### v1

@@ Abstract
Large Reasoning Models (LRMs) perform strongly in complex reasoning tasks via Chain-of-Thought (CoT) prompting, but often suffer from verbose outputs caused by redundant content, increasing computational overhead, and degrading user experience. Existing compression methods either operate post-hoc pruning, risking disruption to reasoning coherence, or rely on sampling-based selection, which fails to intervene effectively during generation. In this work, we introduce a confidence-guided perspective to explain the emergence of redundant reflection in LRMs, identifying two key patterns: Confidence Deficit, where the model reconsiders correct steps due to low internal confidence, and Termination Delay, where reasoning continues even after reaching a confident answer. Based on this analysis, we propose ConCISE (Confidence-guided Compression In Step-by-step Efficient Reasoning), a framework that simplifies reasoning chains by reinforcing the model’s confidence during inference, thus preventing the generation of redundant reflection steps. It integrates Confidence Injection to stabilize intermediate steps and Early Stopping to terminate reasoning when confidence is sufficient. Extensive experiments demonstrate that fine-tuning LRMs on ConCISE-generated data yields significantly shorter outputs, reducing length by up to ~50% under SimPO, while maintaining high task accuracy. ConCISE consistently outperforms existing baselines across multiple reasoning benchmarks.
@@
Ziqing Qiao1, Yongheng Deng1, Jiali Zeng2, Dong Wang1, Lai Wei1, Fandong Meng2, Jie Zhou2, Ju Ren1, Yaoxue Zhang1 1Tsinghua University 2Pattern Recognition Center, WeChat AI, Tencent Inc., China
@@ 1 Introduction
In recent years, Large Language Models (LLMs) have achieved remarkable progress in natural language understanding and generation, particularly in complex reasoning tasks. Large Reasoning Models (LRMs), such as OpenAI-o1 Jaech et al. (2024) and DeepSeek-R1 Guo et al. (2025), which leverage the Chain-of-Thought (CoT) paradigm Wei et al. (2022), have demonstrated state-of-the-art performance on benchmarks involving mathematical problem solving and logical inference Muennighoff et al. (2025); Qwen Team (2024). However, a prominent limitation of LRMs is their tendency to generate excessively verbose reasoning chains Feng et al. (2025); Chen et al. (2024). This inefficiency impairs user experience and incurs considerable computational overhead, posing challenges for real-time applications and deployment in resource-constrained settings Team et al. (2025).
@@
Figure 1: Our proposed method, ConCISE, achieves a more favorable trade-off between reasoning compression and task performance than existing methods.
@@
To mitigate the verbosity of reasoning outputs, recent research has focused on compressing the reasoning chains generated by LRMs Qu et al. (2025); Sui et al. (2025). A mainstream and effective strategy is finetuning LRMs on concise reasoning datasets, enabling the LRMs to generate shorter reasoning chains directly during inference Ma et al. (2025); Chen et al. (2024); Munkhbat et al. (2025). The quality of such compression heavily relies on how these datasets are constructed. One common approach is sampling-based selection, which generates multiple candidate reasoning chains and selects the shortest correct one Team et al. (2025), or removes redundant reflection steps after the final answer is obtained Chen et al. (2024). While simple, these approaches lack control over generation and rely on the reasoning chains naturally generated by the model, which may still include unnecessary steps, reducing compression effectiveness. Another line of work relies on post-hoc pruning, where redundant or less informative reasoning steps are identified and removed from a complete reasoning chain Cui et al. (2025); Xia et al. (2025). However, such modifications can disrupt the continuity of reasoning, leading to degraded performance of LRMs after fine-tuning.
To address the limitations of existing approaches, we seek to construct reasoning chains that are both compact and logically intact, effectively removing redundant reflection steps without sacrificing the performance of LRMs after fine-tuning. To this end, we analyse the reasoning behavior of LRMs and identify that the reflection behavior is not solely determined by correctness, but is heavily influenced by the model’s internal confidence. In particular, insufficient confidence can cause the model to re-evaluate correct steps, resulting in redundant reflections. We thus adopt a confidence-guided perspective to explain why LRMs generate reflection steps. From this view, we identify two key patterns underlying redundant reflection: Confidence Deficit, where the model reconsiders correct intermediate steps due to low internal confidence; and Termination Delay, where reflection continues even after a confident final answer is reached. These patterns substantially inflate reasoning chains and offer actionable insights for constructing concise reasoning data.
Building on this analysis, we propose ConCISE (Confidence-guided Compression In Step-by-step Efficient reasoning)—the first framework that leverages a confidence-guided perspective to actively construct concise reasoning datasets by suppressing redundant reflection during generation. ConCISE consists of two complementary components—Confidence Injection and Early Stopping—designed to mitigate Confidence Deficit and Termination Delay, respectively. Confidence Injection inserts confidence phrases before potential reflection points to strengthen the model’s belief in intermediate steps and suppress unnecessary reflection. Early Stopping employs a lightweight confidence detector to monitor the model’s internal confidence and terminate reasoning once sufficient confidence is achieved. Through the synergy of these two mechanisms, ConCISE enables a more efficient and compact reasoning chain. We then fine-tune LRMs using ConCISE-generated data via SFT Wei et al. (2021) and SimPO Meng et al. (2024), and evaluate their performance on multiple reasoning benchmarks. Experimental results show that ConCISE achieves a superior trade-off between reasoning compression and task performance. As shown in Figure 1, ConCISE reduces average response length by ~50% under SimPO while maintaining high accuracy; under SFT, it also significantly outperforms existing methods in compression efficiency. These results demonstrate that LRMs fine-tuned with ConCISE-generated data learn to avoid generating redundant reflection steps, achieving efficient reasoning without sacrificing performance.
The main contributions of this work are summarized as follows:
  • We present the first confidence-guided explanation of redundant reflection in LRM reasoning, identifying key patterns that underlie reasoning verbosity.
  • We propose ConCISE, a framework that proactively suppresses redundant reflection during inference and constructs high-quality, concise reasoning data for LRM fine-tuning.
  • Through extensive evaluation on multiple reasoning benchmarks, we demonstrate that LRMs fine-tuned with ConCISE-generated data achieve a better trade-off between reasoning chain compression and task performance.
@@ 2 Related Work
Figure 2: The ConCISE reasoning pipeline applied to solving the equation $73x-2=$ 7. The figure compares the original 8-step reasoning process with the streamlined 5-step process generated by ConCISE. It illustrates how ConCISE models reflection based on confidence values, identifying Confidence Deficit and Termination Delay, and applying Confidence Injection and Early Stopping to suppress redundant reasoning. Here, $C_{i}$ denotes the internal confidence of each step, and $T_{i}$ represents its corresponding threshold.
@@
Recent research has increasingly focused on mitigating verbosity and redundancy in reasoning chains generated by Large Reasoning Models (LRMs)Qu et al. (2025). While CoT enables step-by-step reasoning, it often produces long outputs that increase latency, computational costs, and may degrade final accuracyWu et al. (2025); Nayab et al. (2024); Wang et al. (2025). To address this, existing approaches can be broadly categorized into three types: input-based methods, decoding-time interventions, and model-based training strategies Sui et al. (2025).
@@
Input-based Methods aim to promote concise reasoning by modifying input prompts without changing the model architecture or decoding strategyLee et al. (2025). Common techniques include imposing token limits or instructing the model to reason briefly Han et al. (2024); Renze and Guven (2024). However, these approaches often fail when the model does not consistently follow the instructions. Another approach, task routing, dynamically selects an appropriate model or reasoning strategy based on input characteristics like question complexity Ong et al. (2024); Chuang et al. (2025); Aytes et al. (2025). Its effectiveness, however, depends on the accurate assessment of input characteristics, which is not always reliable.
@@
Decoding-based Methods intervene during inference to control the reasoning process. One direction compresses intermediate steps into latent representations rather than explicit text, which improves brevity but sacrifices interpretability Hao et al. (2024); Shen et al. (2025b). Another strategy uses dynamic decoding, where lightweight models or heuristics evaluate each reasoning step and decide whether to retain, modify, or discard it during generation Sun et al. (2024); Zhang et al. (2025); Xu et al. (2025). While effective at reducing verbosity, these methods introduce additional computational overhead and may disrupt the reasoning flow.
@@
Model-based Methods train models to generate concise reasoning directly. One approach employs reinforcement learning with reward functions that penalize verbosity Shen et al. (2025a); Aggarwal and Welleck (2025); Luo et al. (2025); Arora and Zanette (2025), but such methods can be unstable, sensitive to reward formulation, and impose significant training costs. More commonly, LRMs are fine-tuned on datasets with concise reasoning chainsMa et al. (2025); Cui et al. (2025); Liu et al. (2024); Xia et al. (2025). These datasets are often created by selecting the shortest correct chains or removing redundant parts post-hoc. Although effective, these methods lack a principled understanding of redundancy, potentially retaining unnecessary steps or removing useful context, which may hurt compression efficiency or model performance.
@@
Motivated by the limitations of existing methods, we propose a model-based framework ConCISE. Specifically, we achieve precise identification and suppression of redundant reflection steps throughout the entire reasoning process by analyzing and explaining reflection behavior from a confidence-guided perspective. This mechanism allows the model to generate concise and efficient reasoning chains while maintaining the flow of reasoning. By doing so, we construct a high-quality training dataset and effectively transfer this capability to the model through fine-tuning, enabling it to perform concise reasoning without compromising model performance.
@@ 3 Method
3.1 Confidence-guided Formulation of Reflection in LRM Reasoning
We observe that reflection in LRMs is not solely triggered by correctness judgments. In many cases, reflection steps are invoked even on correct reasoning steps, including those that have already been verified, leading to substantial redundancy. This suggests that reflection behavior is also linked to the model’s internal confidence rather than correctness alone. To explain this phenomenon, we adopt a confidence-guided perspective to formalize when and why LRMs engage in reflection.
Let $S_{i}={s_{1},s_{2},\ldots,s_{i}}$ denote the partial reasoning chain up to step $i$, where each $s_{j}$ is a textual reasoning unit. We associate each step $s_{i}$ with a confidence score $c_{i}\in[0,1]$, representing the model’s internal belief in the correctness of that step. The model’s generation policy $\pi_{\theta}$ maps the current reasoning context $S_{i}$ to the next step $s_{i+1}$.
To model the decision between proceeding and reflecting, we introduce a dynamic threshold $t_{i}\in[0,1]$, which may vary with the model architecture or the evolving context. At each step, the model compares its internal confidence $c_{i}$ with this threshold $t_{i}$ to determine whether to continue reasoning or enter reflection. Formally:
$s_{i+1}=\pi_{\theta}(S_{i})=\text{ReflectionStep}\quad\text{if }c_{i}<t_{i}$
Guided by this formulation, we analyze the reflective behavior of LRMs and formally characterize two recurrent patterns responsible for reflection-related redundancy: Confidence Deficit and Termination Delay. These patterns explain how the model’s internal confidence leads to the occurrence of redundant reflection. We will elaborate on each in the following subsection.
(a) Reflection probability after confidence injection.
(b) Detected confidence progression during reasoning.
Figure 3: Validation experiments for Confidence Injection and Termination Delay. The dataset used in these experiments is described in Appendix B.2.
@@ 3.2 Two Key Patterns of Redundancy
One major source of redundancy in LRMs stems from their tendency to undertrust their correct intermediate steps. LRMs often display unexpected hesitation despite exhibiting fine-grained reasoning capabilities and achieving higher stepwise accuracy, triggering reflection even on simple and unambiguous reasoning steps. This self-undermining behavior leads to redundant reflection with minimal semantic gain, ultimately increasing computational cost and disrupting the efficiency of the reasoning process. We refer to this phenomenon as Confidence Deficit. Formally, Confidence Deficit can be defined as the occurrence of a step $s_{i}$ satisfying:
@@
Beyond Confidence Deficit, LRMs exhibit another redundant reasoning pattern we term Termination Delay: After producing a confident final answer, the model is expected to conclude with minimal additional reasoning. However, it often continues to generate unnecessary reflection steps even after repeatedly verifying the same conclusion.
This behavior is attributed to a sharp increase in the confidence threshold $t_{i}$ after reaching the answer at step $i^{\prime}$, making it increasingly difficult for the model’s internal confidence $c_{i}$ to exceed $t_{i}$. As a result, even when $c_{i}$ becomes relatively high, it may still fall short of the heightened $t_{i}$, preventing the model from terminating, leading to verbose post-answer reasoning. Formally, Termination Delay can be characterized as:
@@ 3.3 The ConCISE Framework
To mitigate the above patterns—Confidence Deficit and Termination Delay—we propose ConCISE, an integrated framework that dynamically steers the reasoning generation process to reduce redundant reflections and produce concise reasoning chains.
ConCISE integrates two core mechanisms: Confidence Injection, which proactively inserts confidence phrases to suppress unnecessary reflection steps during reasoning, and Early Stopping, which halts further generation once sufficient post-answer confidence is detected. Together, these mechanisms operate in synergy to generate concise reasoning chains without compromising inference quality. The overall process is illustrated in Figure 2.
@@ 3.3.1 Confidence Injection
To alleviate Confidence Deficit, we design a simple yet effective mechanism called Confidence Injection. The key idea is to boost the model’s internal confidence during reasoning by inserting designed phrases, referred to as confidence phrases, to prevent unnecessary reflection steps. Considering that indiscriminate injection could disrupt reasoning, we selectively insert confidence phrases at critical points where the model’s confidence is relatively low—that is, when $c_{i}<t_{i}$.
Specifically, at each reasoning step $s_{i}$, the model first generates the next step $s_{i+1}=\pi_{\theta}(S_{i})$, where $S_{i}$ denotes the reasoning chain up to step $i$. If $s_{i+1}$ is identified as a reflection step, indicating insufficient confidence, we retroactively modify the input by appending a confidence phrase $p_{i}$ sampled from a curated pool $\mathcal{P}$, and regenerate $s_{i+1}$. Formally, the updated generation process is defined as:
$s_{i+1}=\pi_{\theta}(S_{i})\quad\text{if}\;c_{i}\geq t_{i}\;\text{else}\quad\pi_{\theta}(S_{i},p_{i})$
We manually constructed the initial version of confidence phrase pool $\mathcal{P}$ based on observations of certainty in original reasoning chains. To ensure generalization, the pool includes diverse formulations, and confidence phrases are randomly sampled from it during inference. To evaluate and further refine the confidence phrase pool, we conducted experiments measuring the effectiveness of different phrases. Specifically, we inserted each candidate phrase at reflection positions and recorded the probability that the model still generated a reflection step. This probability indicates whether the updated confidence $c_{i}^{\prime}$ (after phrase insertion) exceeds the dynamic threshold $t_{i}$.
As shown in Figure 3(a), different phrases exhibit varying levels of effectiveness, some notably reduce the likelihood of reflection, while others have only a marginal impact. Notably, even the best-performing phrases still yield a reflection rate above 20%, suggesting that confidence injection reduces redundant reflections while retaining necessary verification. Based on these results, we selected the top 20 phrases with the lowest reflection rates to form the final confidence phrase pool. The composition of the phrase pool and further experimental details are provided in Appendix A.1.

### v2

@@ Abstract
Large Reasoning Models (LRMs) perform strongly in complex reasoning tasks via Chain-of-Thought (CoT) prompting, but often suffer from verbose outputs, increasing computational overhead. Existing fine-tuning-based compression methods either operate post-hoc pruning, risking disruption to reasoning coherence, or rely on sampling-based selection, which fails to remove redundant content thoroughly. To address these limitations, this work begins by framing two key patterns of redundant reflection in LRMs—Confidence Deficit, wherein the model reflects on correct intermediate steps, and Termination Delay, where reflection continues after a verified, confident answer—through a confidence-guided perspective. Based on this, we introduce ConCISE (Confidence-guided Compression In Step-by-step Efficient Reasoning), a framework designed to generate concise reasoning chains, integrating Confidence Injection to boost reasoning confidence, and Early Stopping to terminate reasoning when confidence is sufficient. Extensive experiments demonstrate that compared to baseline methods, fine-tuning LRMs on ConCISE-generated data yields a better balance between compression and task performance, reducing length by up to ~50% under SimPO, while maintaining high task accuracy.
@@
Ziqing Qiao1thanks: Work was done when Ziqing Qiao was interning at Pattern Recognition Center, WeChat AI, Tencent Inc., Yongheng Deng1thanks: Corresponding authors, Jiali Zeng2, Dong Wang1, Lai Wei1, Guanbo Wang1, Fandong Meng [^2] footnotemark: 2, Jie Zhou2, Ju Ren1, Yaoxue Zhang1 1Department of Computer Science and Technology, Tsinghua University 2Pattern Recognition Center, WeChat AI, Tencent Inc., China qzq24@mails.tsinghua.edu.cn dengyh1013@gmail.comfandongmeng@tencent.com
@@ 1 Introduction
In recent years, Large Language Models (LLMs) have achieved remarkable progress in natural language processing, particularly in complex reasoning tasks. Large Reasoning Models (LRMs), such as OpenAI-o1 Jaech et al. (2024) and DeepSeek-R1 Guo et al. (2025), which leverage the Chain-of-Thought paradigm Wei et al. (2022), have demonstrated state-of-the-art performance on benchmarks involving mathematical problem solving and logical reasoningMuennighoff et al. (2025); Qwen Team (2024). However, a prominent limitation of LRMs is their tendency to generate excessively verbose reasoning chains Feng et al. (2025); Chen et al. (2024), incuring considerable computational overhead, and posing challenges for deployment in resource-constrained settings Team et al. (2025).
Figure 1: Training dataset construction workflows: ConCISE (our proposed method) vs. existing methods.
@@
Figure 2: ConCISE achieves a better trade-off between compression and task performance than baselines.
@@
To mitigate LRM output verbosity, recent research focuses on compressing their reasoning chains Qu et al. (2025); Sui et al. (2025). A prominent strategy involves fine-tuning LRMs on concise reasoning datasets, enabling them to generate shorter responses Ma et al. (2025); Chen et al. (2024); Munkhbat et al. (2025). The effectiveness of this compression strategy hinges on the training dataset design. As shown in Figure 1, common sampling-based selection, which generates multiple candidates and picks the shortest correct one, or further removes post-answer redundant reflections Team et al. (2025); Chen et al. (2024), lacks control during generation, potentially leaving unnecessary steps and reducing compression effectiveness. Another approach, post-hoc pruning, identifies and removes redundant or less important steps from reasoning chains Cui et al. (2025); Xia et al. (2025), risks disrupting reasoning coherence and degrading the performance of LRMs after fine-tuning.
To overcome existing limitations, we aim to construct compact, coherent reasoning chains as training datasets by precisely removing redundant reflections, ensuring LRMs do not suffer performance degradation after fine-tuning. To this end, based on the understanding that reflections are not solely determined by correctness Yang et al. (2025), we thus adopt a confidence-guided perspective to understand the generation of reflection steps in LRM’s reasoning processes. This perspective offers a clear interpretation of two key patterns of redundancy: Confidence Deficit, where low internal confidence causes models to undertrust and reflect on correct intermediate steps; and Termination Delay, where reflection persists despite a repeatedly verified answer. These patterns inflate reasoning chains and provide actionable insights for how to create concise reasoning chains.
Therefore, we propose ConCISE (Confidence-guided Compression In Step-by-step Efficient reasoning), a framework that leverages a confidence-guided perspective for constructing concise reasoning data by actively suppressing redundant reflection during generation. ConCISE features two complementary components designed to mitigate Confidence Deficit and Termination Delay, respectively: Confidence Injection, which inserts tailored phrases before potential reflection points to boost the model’s internal confidence, consequently curbing unnecessary reflection; and Early Stopping, which employs a lightweight confidence detector to monitor internal confidence after an answer is reached, terminating reasoning upon achieving sufficient confidence. The synergy of these mechanisms enables ConCISE to produce more efficient and compact reasoning chains.
We then fine-tune four mainstream LRMs using ConCISE-generated data via SFT Wei et al. (2021) and SimPO Meng et al. (2024), evaluating their performance across multiple reasoning benchmarks. As shown in Figure 2, Experimental results demonstrate that ConCISE consistently achieves a superior trade-off between reasoning compression and task performance under both SFT and SimPO settings compared to baselines. Notably, ConCISE reduces average response length by approximately 50% under SimPO while maintaining high accuracy. Our subsequent analysis reveals that ConCISE’s unique training data design enables LRMs, after fine-tuning, to learn to strategically avoid generating redundant reflection steps without harming essential critical reasoning content, thereby achieving efficient reasoning compression while task performance is well maintained.
@@ 2 Related Work
Figure 3: Illustration of ConCISE’s confidence-guided approach: identifying patterns (Confidence Deficit, Termination Delay) and applying mechanisms (Confidence Injection, Early Stopping) to suppress redundant reflections, shown in contrast to the original reasoning process. $C_{i}$ denotes step confidence and $T_{i}$ its threshold.
@@
Recent research has increasingly focused on mitigating verbosity and redundancy in reasoning chains generated by LRMs, which often produces long outputs that increase computational costs, and even degrade accuracy Wu et al. (2025); Nayab et al. (2024); Wang et al. (2025). To address this problem, existing approaches can be broadly categorized into three classes: Input-based, Decoding-based, and Model-based Sui et al. (2025).
@@
Input-based Methods aim to promote concise reasoning by modifying input texts Lee et al. (2025). Common techniques include imposing token limits in prompts or instructing the model to reason briefly Han et al. (2024); Renze and Guven (2024). However, these approaches often fail when the model does not consistently follow these instructions. Another approach, task routing, dynamically selects an appropriate model or reasoning strategy based on input characteristics like question complexity Ong et al. (2024); Chuang et al. (2025); Aytes et al. (2025). Its effectiveness, however, depends on the accurate assessment of input characteristics, which is not always reliable.
@@
Decoding-based Methods intervene during decoding to control the reasoning process. One common approach compresses steps into latent representations rather than explicit text, which improves brevity but sacrifices interpretability Hao et al. (2024); Shen et al. (2025b). Another strategy uses dynamic decoding, evaluating each reasoning step and deciding whether to modify or discard it during generation Sun et al. (2024); Zhang et al. (2025); Xu et al. (2025). While effective at reducing verbosity, these methods introduce additional computational overhead and may disrupt the reasoning coherence.
@@
Model-based Methods train models to generate concise reasoning directly. One approach employs reinforcement learning with reward functions that penalize verbosity Shen et al. (2025a); Aggarwal and Welleck (2025); Luo et al. (2025); Arora and Zanette (2025); Yu et al. (2025), but can be sensitive to reward formulation and impose significant costs. Another common approach is to fine-tune LRMs on datasets with concise reasoning chainsMa et al. (2025); Cui et al. (2025); Xia et al. (2025). These datasets are often created by selecting the shortest correct chain or removing redundant parts post-hoc. Such methods may inadvertently retain redundant reflection steps or remove useful context, degrading compression efficiency or model performance.
@@
Motivated by the limitations of existing methods, we propose ConCISE, a model-based method that precisely identifies and actively suppresses redundant reflection steps throughout the reasoning process. This approach enables the model to generate efficient reasoning chains while preserving reasoning coherence. Leveraging ConCISE, we construct a high-quality training dataset and fine-tune the model to adopt this concise reasoning ability without compromising model performance.
@@ 3 Method
3.1 Confidence-guided Formulation
Reflections in LRMs are not solely triggered by correctness; in many cases, reflection steps are invoked even on correct steps that have been verified Yang et al. (2025). This suggests that reflection behavior is also linked to the model’s internal confidence about current reasoning rather than correctness alone. To explain this, we adopt a confidence-guided perspective to formalize when and why LRMs engage in reflection.
Let $S_{i}=\{s_{1},s_{2},\ldots,s_{i}\}$ denote the partial reasoning chain up to step $i$, where each $s_{i}$ is a textual reasoning unit. We associate each step $s_{i}$ with a confidence score $c_{i}\in[0,1]$, representing the model’s internal confidence of that step. The generation policy of the LRM, denoted by $\pi_{\theta}$, maps the current reasoning context $S_{i}$ to the next step $s_{i+1}$.
To model the decision between proceeding and reflecting, we introduce a dynamic threshold $t_{i}\in[0,1]$, which may vary with the model or the context. At each step, the model generates a ReflectionStep as $s_{i+1}$ if its current internal confidence $c_{i}$ falls below the threshold $t_{i}$ (i.e., $c_{i}<t_{i}$). From this perspective, we further analyze the reflection behavior of LRMs and formally articulate two key patterns responsible for reflection-related redundancy existing in the reasoning process: Confidence Deficit and Termination Delay.
@@ 3.2 Two Key Patterns of Redundancy
One major source of redundancy in LRMs stems from their tendency to undertrust their correct intermediate steps. LRMs often display unexpected reflection despite exhibiting fine-grained reasoning capabilities and achieving high stepwise accuracy, triggering reflection even on simple and unambiguous reasoning steps. This self-undermining behavior leads to redundant reflection with minimal semantic gain. We refer to this phenomenon as Confidence Deficit. Formally, Confidence Deficit can be defined as the occurrence of a step $s_{i}$ satisfying:
@@
LRMs exhibit another important redundant reasoning pattern we term Termination Delay: After producing a confident final answer, the model is expected to conclude with minimal additional reasoning. However, it often continues to generate unnecessary reflection steps even after repeatedly verifying the same conclusion. This behavior can be attributed to a sharp increase in the confidence threshold $t_{i}$ after reaching the answer at step $i^{\prime}$, making it increasingly difficult for the model’s internal confidence $c_{i}$ to exceed $t_{i}$. As a result, even when $c_{i}$ becomes relatively high, it may still fall short of the heightened $t_{i}$, leading to verbose post-answer reasoning. Formally, Termination Delay can be characterized as:
@@ 3.3 The ConCISE Framework
To mitigate the above patterns—Confidence Deficit and Termination Delay—we propose ConCISE, a framework that dynamically steers the reasoning generation process to reduce redundant reflections and produce concise reasoning chains. ConCISE integrates two mechanisms: Confidence Injection, which actively inserts confidence phrases to suppress unnecessary reflection steps, and Early Stopping, which halts generation once sufficient post-answer confidence is detected. These two mechanisms operate together to generate concise reasoning chains without compromising inference quality. The overall process is illustrated in Figure 3.
@@ 3.3.1 Confidence Injection
To alleviate Confidence Deficit, we design a simple yet effective mechanism called Confidence Injection. The key idea is to boost the model’s internal confidence during reasoning by inserting designed phrases, referred to as confidence phrases, to prevent unnecessary reflection steps. Considering that indiscriminate injection could disrupt reasoning, we selectively insert confidence phrases at critical points where the model’s confidence is relatively low, that is, when $c_{i}<t_{i}$, a condition that signifies an impending reflection step.
Specifically, at each step $s_{i}$, the model first generates $s_{i+1}=\pi_{\theta}(S_{i})$. If $s_{i+1}$ is a reflection step, we retroactively modify the input $S_{i}$ by appending a confidence phrase $p_{i}$ sampled from a curated pool $\mathcal{P}$, and regenerate $s_{i+1}$. Formally, the updated generation process is defined as:
$s_{i+1}=\pi_{\theta}(S_{i})\quad\text{if}\;c_{i}\geq t_{i}\;\text{else}\;\pi_{\theta}(S_{i},p_{i})$
We manually constructed the initial version of $\mathcal{P}$. To evaluate and further refine the initial pool, we conducted experiments measuring the effectiveness of different phrases. We inserted each candidate phrase before the reflection steps and recorded the probability that the model still generated a reflection step. This probability indicates whether the updated confidence $c_{i}^{*}$ (after phrase insertion) exceeds the dynamic threshold $t_{i}$. As shown in Figure 4(a), different phrases exhibit varying reflection rates. Notably, even the best-performing phrases still yield a reflection rate around 20%, suggesting that confidence injection reduces redundant reflections while still retaining necessary verifications. We then selected 20 phrases with the lowest reflection rates to form the final confidence phrase pool. The composition of the phrase pool and detection of reflection steps are provided in Appendix A.1.